# Chain-of-Thought in Neural Code Generation: From and For Lightweight Language Models

### v1

@@ 5.2 RQ2: Can COTTON generate higher-quality CoTs
Self-CoT CodeLlama ChatGLM(130B) Teacher(175B) COTTON
HE CodeGen 350M 14.63 12.80 18.90 ($\uparrow$ 29.19%) 18.29 ($\uparrow$ 25.02%) 24.39 ($\uparrow$ 66.71%) 20.73 ($\uparrow$ 41.70%)
CodeGen 2B 25.61 25.00 31.10 ($\uparrow$ 21.44%) 31.71 ($\uparrow$ 23.82%) 41.46 ($\uparrow$ 61.89%) 34.76 ($\uparrow$ 35.73%)
CodeGen 6B 27.44 31.10 ($\uparrow$ 13.34%) 36.59 ($\uparrow$ 33.35%) 31.71 ($\uparrow$ 15.56%) 43.29 ($\uparrow$ 57.76%) 39.63 ($\uparrow$ 44.42%)
StarCoder 1B 12.80 12.80 17.68 ($\uparrow$ 38.13%) 17.68 ($\uparrow$ 38.13%) 28.66 ($\uparrow$ 123.91%) 25.00 ($\uparrow$ 95.31%)
StarCoder 3B 17.07 21.34 ($\uparrow$ 25.01%) 25.61 ($\uparrow$ 50.03%) 25.61 ($\uparrow$ 50.03%) 39.63 ($\uparrow$ 132.16%) 30.49 ($\uparrow$ 78.62%)
StarCoder 7B 21.95 29.88 ($\uparrow$ 36.13%) 33.54 ($\uparrow$ 52.80%) 34.15 ($\uparrow$ 35.72%) 41.46 ($\uparrow$ 88.88%) 37.20 ($\uparrow$ 69.48%)
CodeT5+ 220M 12.20 14.02 ($\uparrow$ 14.92%) 18.29 ($\uparrow$ 49.92%) 16.46 ($\uparrow$ 34.92%) 23.17 ($\uparrow$ 89.92%) 18.90 ($\uparrow$ 54.92%)
CodeT5+ 770M 17.07 17.68 ($\uparrow$ 3.57%) 23.78 ($\uparrow$ 39.31%) 23.78 ($\uparrow$ 39.31%) 31.71 ($\uparrow$ 85.76%) 26.83 ($\uparrow$ 57.18%)
CodeT5+ 2B 23.78 25.00 ($\uparrow$ 5.13%) 28.05 ($\uparrow$ 17.96%) 29.27 ($\uparrow$ 23.09%) 38.41 ($\uparrow$ 61.52%) 30.49 ($\uparrow$ 28.22%)
CodeT5+ 6B 26.22 32.32 ($\uparrow$ 23.26%) 38.41 ($\uparrow$ 46.69%) 36.59 ($\uparrow$ 39.55%) 47.56 ($\uparrow$ 81.39%) 42.68 ($\uparrow$ 62.78%)
OE CodeGen 350M 7.30 8.99 ($\uparrow$ 23.15%) 14.04 ($\uparrow$ 92.33%) 12.92 ($\uparrow$ 76.99%) 15.17 ($\uparrow$ 107.81%) 12.92 ($\uparrow$ 76.99%)
CodeGen 2B 16.85 19.10 ($\uparrow$ 13.35%) 21.91 ($\uparrow$ 30.03%) 25.28 ($\uparrow$ 50.03%) 29.21 ($\uparrow$ 73.35%) 26.97 ($\uparrow$ 60.06%)
CodeGen 6B 21.91 23.60 ($\uparrow$ 7.71%) 29.21 ($\uparrow$ 33.32%) 32.02 ($\uparrow$ 46.14%) 37.64 ($\uparrow$ 71.79%) 33.71 ($\uparrow$ 53.86%)
StarCoder 1B 8.99 10.67 ($\uparrow$ 18.69%) 14.61 ($\uparrow$ 62.51%) 16.85 ($\uparrow$ 87.43%) 19.66 ($\uparrow$ 118.69%) 17.42 ($\uparrow$ 93.77%)
StarCoder 3B 11.24 16.85 ($\uparrow$ 49.91%) 15.17 ($\uparrow$ 34.96%) 14.61 ($\uparrow$ 29.98%) 17.98 ($\uparrow$ 59.96%) 19.10 ($\uparrow$ 69.93%)
StarCoder 7B 23.03 29.21 ($\uparrow$ 26.83%) 28.65 ($\uparrow$ 24.40%) 29.78 ($\uparrow$ 29.31%) 38.20 ($\uparrow$ 65.87%) 33.15 ($\uparrow$ 43.94%)
CodeT5+ 220M 7.87 10.67 ($\uparrow$ 35.58%) 10.67 ($\uparrow$ 35.58%) 12.92 ($\uparrow$ 64.17%) 14.04 ($\uparrow$ 78.40%) 13.48 ($\uparrow$ 71.28%)
CodeT5+ 770M 9.55 11.24 ($\uparrow$ 17.70%) 16.85 ($\uparrow$ 76.44%) 17.98 ($\uparrow$ 88.27%) 20.22 ($\uparrow$ 111.73%) 16.29 ($\uparrow$ 70.58%)
CodeT5+ 2B 15.17 19.66 ($\uparrow$ 29.60%) 24.72 ($\uparrow$ 62.95%) 25.28 ($\uparrow$ 66.64%) 30.90 ($\uparrow$ 103.69%) 28.65 ($\uparrow$ 88.86%)
CodeT5+ 6B 20.22 21.35 ($\uparrow$ 5.59%) 28.09 ($\uparrow$ 38.92%) 32.02 ($\uparrow$ 58.36%) 36.52 ($\uparrow$ 80.61%) 35.39 ($\uparrow$ 75.02%)
HE-p CodeGen 350M 15.24 12.80 19.51 ($\uparrow$ 28.02%) 18.29 ($\uparrow$ 20.01%) 25.00 ($\uparrow$ 64.04%) 20.73 ($\uparrow$ 36.02%)
CodeGen 2B 26.22 25.00 31.71 ($\uparrow$ 20.94%) 32.32 ($\uparrow$ 23.26%) 42.68 ($\uparrow$ 62.78%) 35.37 ($\uparrow$ 34.90%)
CodeGen 6B 27.44 32.32 ($\uparrow$ 17.78%) 37.20 ($\uparrow$ 35.57%) 32.32 ($\uparrow$ 17.78%) 44.51 ($\uparrow$ 62.21%) 40.85 ($\uparrow$ 48.87%)
StarCoder 1B 13.41 13.41 18.29 ($\uparrow$ 36.39%) 18.29 ($\uparrow$ 36.39%) 29.27 ($\uparrow$ 118.27%) 26.22 ($\uparrow$ 95.53%)
StarCoder 3B 17.07 21.95 ($\uparrow$ 28.59%) 25.61 ($\uparrow$ 50.03%) 26.22 ($\uparrow$ 53.60%) 40.85 ($\uparrow$ 139.31%) 31.71 ($\uparrow$ 85.76%)
StarCoder 7B 22.56 30.49 ($\uparrow$ 35.15%) 34.15 ($\uparrow$ 51.37%) 34.76 ($\uparrow$ 54.08%) 43.29 ($\uparrow$ 91.89%) 38.41 ($\uparrow$ 70.26%)
CodeT5+ 220M 12.20 14.02 ($\uparrow$ 14.92%) 18.90 ($\uparrow$ 54.92%) 17.07 ($\uparrow$ 39.92%) 23.78 ($\uparrow$ 94.92%) 19.51 ($\uparrow$ 59.92%)
CodeT5+ 770M 17.68 17.68 24.39 ($\uparrow$ 37.95%) 24.39 ($\uparrow$ 37.95%) 32.32 ($\uparrow$ 82.81%) 27.44 ($\uparrow$ 55.20%)
CodeT5+ 2B 25.00 25.00 28.66 ($\uparrow$ 14.64%) 29.88 ($\uparrow$ 19.52%) 39.63 ($\uparrow$ 58.52%) 31.71 ($\uparrow$ 26.84%)
CodeT5+ 6B 26.83 32.93 ($\uparrow$ 22.74%) 39.02 ($\uparrow$ 45.43%) 37.20 ($\uparrow$ 38.65%) 48.78 ($\uparrow$ 81.81%) 43.90 ($\uparrow$ 63.62%)
@@
TABLE VII: The performance of code generation models with or without the CoTs generated by COTTON using different base models on the HumanEval(HE), HumanEval-plus(HE-p), and OpenEval(OE) datasets
@@
GraphCodeBERT CodeGPT-adapter NatGen CodeGeeX2 CodeLlama

### v2

@@ 5.2 RQ2: Can COTTON generate higher-quality CoTs
Self-CoT Think Step by Step Self-planning SCoT COTTON
HE CodeGen 350M 14.63 12.80 15.85 ($\uparrow$ 8.34%) 18.29 ($\uparrow$ 25.02%) 15.85 ($\uparrow$ 8.34%) 20.73 ($\uparrow$ 41.70%)
CodeGen 2B 25.61 25.00 27.44 ($\uparrow$ 7.15%) 26.22 ($\uparrow$ 2.38%) 26.83 ($\uparrow$ 4.76%) 34.76 ($\uparrow$ 35.73%)
CodeGen 6B 27.44 31.10 ($\uparrow$ 13.34%) 29.88 ($\uparrow$ 8.89%) 28.66 ($\uparrow$ 4.45%) 30.49 ($\uparrow$ 11.12%) 39.63 ($\uparrow$ 44.42%)
StarCoder 1B 12.80 12.80 15.85 ($\uparrow$ 23.83%) 17.68 ($\uparrow$ 38.13%) 14.63 ($\uparrow$ 14.30%) 25.00 ($\uparrow$ 95.31%)
StarCoder 3B 17.07 21.34 ($\uparrow$ 25.01%) 17.07 20.12 ($\uparrow$ 17.87%) 20.12 ($\uparrow$ 17.87%) 30.49 ($\uparrow$ 78.62%)
StarCoder 7B 21.95 29.88 ($\uparrow$ 36.13%) 28.05 ($\uparrow$ 27.79%) 25.61 ($\uparrow$ 16.67%) 29.88 ($\uparrow$ 36.13%) 37.20 ($\uparrow$ 69.48%)
CodeT5+ 220M 12.20 14.02 ($\uparrow$ 14.92%) 14.02 ($\uparrow$ 14.92%) 12.20 11.59 18.90 ($\uparrow$ 54.92%)
CodeT5+ 770M 17.07 17.68 ($\uparrow$ 3.57%) 17.68 ($\uparrow$ 3.57%) 17.68 ($\uparrow$ 3.57%) 18.90 ($\uparrow$ 10.72%) 26.83 ($\uparrow$ 57.18%)
CodeT5+ 2B 23.78 25.00 ($\uparrow$ 5.13%) 27.44 ($\uparrow$ 15.39%) 26.22 ($\uparrow$ 10.26%) 26.22 ($\uparrow$ 10.26%) 30.49 ($\uparrow$ 28.22%)
CodeT5+ 6B 26.22 32.32 ($\uparrow$ 23.26%) 33.54 ($\uparrow$ 27.92%) 34.15 ($\uparrow$ 30.24%) 32.93 ($\uparrow$ 25.59%) 42.68 ($\uparrow$ 62.78%)
OE CodeGen 350M 7.30 8.99 ($\uparrow$ 23.15%) 8.43 ($\uparrow$ 15.48%) 7.87 ($\uparrow$ 7.81%) 10.67 ($\uparrow$ 46.16%) 12.92 ($\uparrow$ 76.99%)
CodeGen 2B 16.85 19.10 ($\uparrow$ 13.35%) 16.85 20.22 ($\uparrow$ 20.00%) 20.22 ($\uparrow$ 20.00%) 26.97 ($\uparrow$ 60.06%)
CodeGen 6B 21.91 23.60 ($\uparrow$ 7.71%) 23.60 ($\uparrow$ 7.71%) 28.09 ($\uparrow$ 28.21%) 29.78 ($\uparrow$ 35.92%) 33.71 ($\uparrow$ 53.86%)
StarCoder 1B 8.99 10.67 ($\uparrow$ 18.69%) 8.99 11.24 ($\uparrow$ 25.03%) 11.80 ($\uparrow$ 31.26%) 17.42 ($\uparrow$ 93.77%)
StarCoder 3B 11.24 16.85 ($\uparrow$ 49.91%) 14.61 ($\uparrow$ 29.98%) 19.10 ($\uparrow$ 69.93%) 14.61 ($\uparrow$ 29.98%) 19.10 ($\uparrow$ 69.93%)
StarCoder 7B 23.03 29.21 ($\uparrow$ 26.83%) 28.65 ($\uparrow$ 24.40%) 25.84 ($\uparrow$ 12.20%) 28.09 ($\uparrow$ 21.97%) 33.15 ($\uparrow$ 43.94%)
CodeT5+ 220M 7.87 10.67 ($\uparrow$ 35.58%) 12.36 ($\uparrow$ 57.05%) 10.67 ($\uparrow$ 35.58%) 8.43 ($\uparrow$ 7.12%) 13.48 ($\uparrow$ 71.28%)
CodeT5+ 770M 9.55 11.24 ($\uparrow$ 17.70%) 13.48 ($\uparrow$ 41.15%) 13.48 ($\uparrow$ 41.15%) 13.48 ($\uparrow$ 41.15%) 16.29 ($\uparrow$ 70.58%)
CodeT5+ 2B 15.17 19.66 ($\uparrow$ 29.60%) 16.85 ($\uparrow$ 11.07%) 19.66 ($\uparrow$ 29.60%) 17.42 ($\uparrow$ 14.83%) 28.65 ($\uparrow$ 88.86%)
CodeT5+ 6B 20.22 21.35 ($\uparrow$ 5.59%) 30.90 ($\uparrow$ 52.82%) 24.72 ($\uparrow$ 22.26%) 31.46 ($\uparrow$ 55.59%) 35.39 ($\uparrow$ 75.02%)
HE-p CodeGen 350M 15.24 12.80 16.46 ($\uparrow$ 8.01%) 18.29 ($\uparrow$ 20.01%) 15.85 ($\uparrow$ 4.00%) 20.73 ($\uparrow$ 36.02%)
CodeGen 2B 26.22 25.00 27.44 ($\uparrow$ 4.65%) 26.22 26.83 ($\uparrow$ 2.33%) 35.37 ($\uparrow$ 34.90%)
CodeGen 6B 27.44 32.32 ($\uparrow$ 17.78%) 30.49 ($\uparrow$ 11.12%) 29.27 ($\uparrow$ 6.67%) 30.49 ($\uparrow$ 11.12%) 40.85 ($\uparrow$ 48.87%)
StarCoder 1B 13.41 13.41 15.85 ($\uparrow$ 18.20%) 17.68 ($\uparrow$ 31.84%) 14.63 ($\uparrow$ 9.10%) 26.22 ($\uparrow$ 95.53%)
StarCoder 3B 17.07 21.95 ($\uparrow$ 28.59%) 17.07 20.12 ($\uparrow$ 17.87%) 20.73 ($\uparrow$ 21.44%) 31.71 ($\uparrow$ 85.76%)
StarCoder 7B 22.56 30.49 ($\uparrow$ 35.15%) 28.66 ($\uparrow$ 27.04%) 25.61 ($\uparrow$ 13.52%) 31.10 ($\uparrow$ 37.85%) 38.41 ($\uparrow$ 70.26%)
CodeT5+ 220M 12.20 14.02 ($\uparrow$ 14.92%) 14.02 ($\uparrow$ 14.92%) 12.20 11.59 19.51 ($\uparrow$ 59.92%)
CodeT5+ 770M 17.68 17.68 17.68 17.68 19.51 ($\uparrow$ 10.35%) 27.44 ($\uparrow$ 55.20%)
CodeT5+ 2B 25.00 25.00 28.05 ($\uparrow$ 12.20%) 27.44 ($\uparrow$ 9.76%) 26.22 ($\uparrow$ 4.88%) 31.71 ($\uparrow$ 26.84%)
CodeT5+ 6B 26.83 32.93 ($\uparrow$ 22.74%) 34.15 ($\uparrow$ 27.28%) 34.76 ($\uparrow$ 26.56%) 32.93 ($\uparrow$ 22.74%) 43.90 ($\uparrow$ 63.62%)
For the human evaluation, we recruited six assessors (six postgraduate students) who were familiar with Python programming language. We selected all samples from the HumanEval-CoT dataset (164 samples) and the OpenEval-CoT dataset (178 samples) as the evaluation subjects. For each sample, we collected the ground truth CoT and three generated CoTs.
To ensure a thorough evaluation, we divided all the samples into three groups, with each group containing 114 samples. Each group was evaluated anonymously by two assessors in terms of similarity, naturalness, and educational value. The score for each aspect ranges from 0 to 4, with higher scores indicating higher quality and the final score is the average of two assessors’ scores.
A sample questionnaire is shown in Figure 4. To guarantee human study quality, the generated CoTs were presented in a random order, ensuring that the assessors had no knowledge of which approach generated the CoT. Moreover, the assessors were allowed to use the internet to look up any related concepts they were unfamiliar with. Finally, we limited each assessor to evaluate only 20 samples in half a day. This was done to prevent fatigue and maintain a high level of concentration during the evaluation process.
The results of the human study are summarized in Table V, which shows the average scores and standard deviations of all evaluated samples by the assessors. We can observe that the proposed approach COTTON outperforms Llama2 and CodeGeeX2 in terms of similarity, naturalness, and educational value. To determine the statistical significance of these differences, we conducted Fisher’s exact test [64] and observed a statistically significant disparity in Table VI (i.e., $p$-value $<0.05$). Furthermore, we utilized Fleiss Kappa [65] to assess the agreement among the six assessors. The overall Kappa value based on the comparison results is 0.7, signifying substantial agreement among the assessors.
These findings provide further evidence of the effectiveness of COTTON in generating high-quality CoTs for guiding code generation. The higher average score values obtained by COTTON indicate that it is capable of generating CoTs that are more grammatically correct, fluent, and valuable in terms of providing guidance and inspiration to developers. These results highlight the competitiveness of our proposed approach in generating CoTs that are not only grammatically sound but also valuable from an educational perspective.
Summary of RQ2 COTTON consistently outperforms the state-of-the-art base models in generating CoTs in terms of lexical and semantic similarity, as well as its ability to provide valuable guidance and inspiration to developers.
@@
TABLE VIII: The performance of code generation models with or without the CoTs generated by different methods on the HumanEval(HE), HumanEval-plus(HE-p), and OpenEval(OE) datasets (Here COTTON is based on CodeLlama-7b).
Corpus Model Pass@1 CoT-Pass@1 (different CoT generation methods)
Gemini CodeLlama ChatGLM Teacher COTTON
HE CodeGen 350M 14.63 24.39 ($\uparrow$ 66.71%) 18.90 ($\uparrow$ 29.19%) 18.29 ($\uparrow$ 25.02%) 24.39 ($\uparrow$ 66.71%) 20.73 ($\uparrow$ 41.70%)
CodeGen 2B 25.61 37.20 ($\uparrow$ 45.26%) 31.10 ($\uparrow$ 21.44%) 31.71 ($\uparrow$ 23.82%) 41.46 ($\uparrow$ 61.89%) 34.76 ($\uparrow$ 35.73%)
CodeGen 6B 27.44 43.29 ($\uparrow$ 57.76%) 36.59 ($\uparrow$ 33.35%) 31.71 ($\uparrow$ 15.56%) 43.29 ($\uparrow$ 57.76%) 39.63 ($\uparrow$ 44.42%)
StarCoder 1B 12.80 21.95 ($\uparrow$ 71.48%) 17.68 ($\uparrow$ 38.13%) 17.68 ($\uparrow$ 38.13%) 28.66 ($\uparrow$ 123.91%) 25.00 ($\uparrow$ 95.31%)
StarCoder 3B 17.07 34.15 ($\uparrow$ 100.06%) 25.61 ($\uparrow$ 50.03%) 25.61 ($\uparrow$ 50.03%) 39.63 ($\uparrow$ 132.16%) 30.49 ($\uparrow$ 78.62%)
StarCoder 7B 21.95 38.41 ($\uparrow$ 74.99%) 33.54 ($\uparrow$ 52.80%) 34.15 ($\uparrow$ 35.72%) 41.46 ($\uparrow$ 88.88%) 37.20 ($\uparrow$ 69.48%)
CodeT5+ 220M 12.20 19.51 ($\uparrow$ 59.92%) 18.29 ($\uparrow$ 49.92%) 16.46 ($\uparrow$ 34.92%) 23.17 ($\uparrow$ 89.92%) 18.90 ($\uparrow$ 54.92%)
CodeT5+ 770M 17.07 26.83 ($\uparrow$ 57.18%) 23.78 ($\uparrow$ 39.31%) 23.78 ($\uparrow$ 39.31%) 31.71 ($\uparrow$ 85.76%) 26.83 ($\uparrow$ 57.18%)
CodeT5+ 2B 23.78 38.41 ($\uparrow$ 61.52%) 28.05 ($\uparrow$ 17.96%) 29.27 ($\uparrow$ 23.09%) 38.41 ($\uparrow$ 61.52%) 30.49 ($\uparrow$ 28.22%)
CodeT5+ 6B 26.22 45.73 ($\uparrow$ 74.41%) 38.41 ($\uparrow$ 46.69%) 36.59 ($\uparrow$ 39.55%) 47.56 ($\uparrow$ 81.39%) 42.68 ($\uparrow$ 62.78%)
OE CodeGen 350M 7.30 16.85 ($\uparrow$ 13.08%) 14.04 ($\uparrow$ 92.33%) 12.92 ($\uparrow$ 76.99%) 15.17 ($\uparrow$ 107.81%) 12.92 ($\uparrow$ 76.99%)
CodeGen 2B 16.85 30.34 ($\uparrow$ 80.06%) 21.91 ($\uparrow$ 30.03%) 25.28 ($\uparrow$ 50.03%) 29.21 ($\uparrow$ 73.35%) 26.97 ($\uparrow$ 60.06%)
CodeGen 6B 21.91 34.27 ($\uparrow$ 56.41%) 29.21 ($\uparrow$ 33.32%) 32.02 ($\uparrow$ 46.14%) 37.64 ($\uparrow$ 71.79%) 33.71 ($\uparrow$ 53.86%)
StarCoder 1B 8.99 17.42 ($\uparrow$ 93.77%) 14.61 ($\uparrow$ 62.51%) 16.85 ($\uparrow$ 87.43%) 19.66 ($\uparrow$ 118.69%) 17.42 ($\uparrow$ 93.77%)
StarCoder 3B 11.24 18.54 ($\uparrow$ 64.95%) 15.17 ($\uparrow$ 34.96%) 14.61 ($\uparrow$ 29.98%) 17.98 ($\uparrow$ 59.96%) 19.10 ($\uparrow$ 69.93%)
StarCoder 7B 23.03 34.83 ($\uparrow$ 51.24%) 28.65 ($\uparrow$ 24.40%) 29.78 ($\uparrow$ 29.31%) 38.20 ($\uparrow$ 65.87%) 33.15 ($\uparrow$ 43.94%)
CodeT5+ 220M 7.87 15.17 ($\uparrow$ 92.76%) 10.67 ($\uparrow$ 35.58%) 12.92 ($\uparrow$ 64.17%) 14.04 ($\uparrow$ 78.40%) 13.48 ($\uparrow$ 71.28%)
CodeT5+ 770M 9.55 21.35 ($\uparrow$ 123.56%) 16.85 ($\uparrow$ 76.44%) 17.98 ($\uparrow$ 88.27%) 20.22 ($\uparrow$ 111.73%) 16.29 ($\uparrow$ 70.58%)
CodeT5+ 2B 15.17 28.65 ($\uparrow$ 88.86%) 24.72 ($\uparrow$ 62.95%) 25.28 ($\uparrow$ 66.64%) 30.90 ($\uparrow$ 103.69%) 28.65 ($\uparrow$ 88.86%)
CodeT5+ 6B 20.22 33.15 ($\uparrow$ 63.95%) 28.09 ($\uparrow$ 38.92%) 32.02 ($\uparrow$ 58.36%) 36.52 ($\uparrow$ 80.61%) 35.39 ($\uparrow$ 75.02%)
HE-p CodeGen 350M 15.24 24.39 ($\uparrow$ 60.04%) 19.51 ($\uparrow$ 28.02%) 18.29 ($\uparrow$ 20.01%) 25.00 ($\uparrow$ 64.04%) 20.73 ($\uparrow$ 36.02%)
CodeGen 2B 26.22 37.80 ($\uparrow$ 44.16%) 31.71 ($\uparrow$ 20.94%) 32.32 ($\uparrow$ 23.26%) 42.68 ($\uparrow$ 62.78%) 35.37 ($\uparrow$ 34.90%)
CodeGen 6B 27.44 43.90 ($\uparrow$ 59.99%) 37.20 ($\uparrow$ 35.57%) 32.32 ($\uparrow$ 17.78%) 44.51 ($\uparrow$ 62.21%) 40.85 ($\uparrow$ 48.87%)
StarCoder 1B 13.41 22.56 ($\uparrow$ 68.23%) 18.29 ($\uparrow$ 36.39%) 18.29 ($\uparrow$ 36.39%) 29.27 ($\uparrow$ 118.27%) 26.22 ($\uparrow$ 95.53%)
StarCoder 3B 17.07 34.76 ($\uparrow$ 103.63%) 25.61 ($\uparrow$ 50.03%) 26.22 ($\uparrow$ 53.60%) 40.85 ($\uparrow$ 139.31%) 31.71 ($\uparrow$ 85.76%)
StarCoder 7B 22.56 39.02 ($\uparrow$ 72.96%) 34.15 ($\uparrow$ 51.37%) 34.76 ($\uparrow$ 54.08%) 43.29 ($\uparrow$ 91.89%) 38.41 ($\uparrow$ 70.26%)
CodeT5+ 220M 12.20 20.12 ($\uparrow$ 64.92%) 18.90 ($\uparrow$ 54.92%) 17.07 ($\uparrow$ 39.92%) 23.78 ($\uparrow$ 94.92%) 19.51 ($\uparrow$ 59.92%)
CodeT5+ 770M 17.68 27.44 ($\uparrow$ 55.20%) 24.39 ($\uparrow$ 37.95%) 24.39 ($\uparrow$ 37.95%) 32.32 ($\uparrow$ 82.81%) 27.44 ($\uparrow$ 55.20%)
CodeT5+ 2B 25.00 39.02 ($\uparrow$ 56.08%) 28.66 ($\uparrow$ 14.64%) 29.88 ($\uparrow$ 19.52%) 39.63 ($\uparrow$ 58.52%) 31.71 ($\uparrow$ 26.84%)
CodeT5+ 6B 26.83 46.34 ($\uparrow$ 72.72%) 39.02 ($\uparrow$ 45.43%) 37.20 ($\uparrow$ 38.65%) 48.78 ($\uparrow$ 81.81%) 43.90 ($\uparrow$ 63.62%)
TABLE IX: The performance of code generation models with or without the CoTs generated by COTTON using different base models on the HumanEval(HE), HumanEval-plus(HE-p), and OpenEval(OE) datasets (Here COTTON is based on CodeLlama-7b).
@@
GraphCodeBERT CodeGPT-adapter NatGen CodeGeeX2 COTTON